A cat leaps trying to catch a flye. The motion of the cat can be modeled by h(d)=-32d^2+16d. The fly is staying at least 2.5 feet away from the ground. Will the cat get high enough to catch the fly?
Answer by stanbon(75887)

A cat leaps trying to catch a fly. The motion of the cat can be modeled by
h(d)=-32d^2+16d. The fly is staying at least 2.5 feet away from the ground. Will the cat get high enough to catch the fly?
-----------------
h(d)=-32d^2+16d
---
Maximum height occurs when d = -b/2a = -16/(2*-32) = 1/4
---
That maximum height is h(1/4) = -32(1/4)^2+16(1/4) = 2 ft.
----
I guess the cat is going to go hungry.
============================
Cheers,
Stan H.